# Synomaly Noise and Multi-Stage Diffusion: A Novel Approach for Unsupervised Anomaly Detection in Ultrasound Imaging

### v1

@@ 4.7.2 Ablation for Multi-stage Diffusion Process
To validate the effectiveness of the proposed multi-stage diffusion process, we conduct an ablation experiment to compare the anomaly segmentation performance with and without multi-stage diffusion. The DDPM was trained using Synomaly noise, and during inference, we tested the model’s performance with a single-stage diffusion approach. To ensure a fair comparison, the optimal inference hyperparameters for single-stage diffusion with Synomaly noise were determined using a grid search and are reported in Tab. 1, denoted as “Synomaly+Single-stage”. In general, the number of noise steps at each stage for multi-stage inference is much smaller than the number of noise steps for single stage, which helps to preserve finer details. The quantitative results can be found in Tab. 3 under the “Synomaly+Single-stage”, and the visual results can be found in Fig. 7 on the right part.
@@
For US carotid plaque anomaly segmentation, implementing the multi-stage diffusion process has improved the Dice score by approximately compared to the single-stage approach. The two visual examples in Fig. 7(g) and Fig. 7(h) also demonstrate that a single diffusion process cannot fully erase the anomalous regions. While the plaque in the upper part of the vessel is successfully removed, the larger plaque at the bottom of the vessel remains in the generated image. After applying multi-stage diffusion process, all the plaques are fully removed. Fig. 7(c) shows the generated images at each stage, clearly illustrating how the plaques are iteratively removed from the original image. Generally, distinct and large anomalies require more iterations to be completely erased. Fig. 7(d) showcases another example involving five iterations, where the distinct plaque at the bottom of the vessel lumen is progressively removed over five iterations, while the smaller plaque at the top is mostly removed in four iterations.
@@
5 Conclusion and Discussion
@@
In this paper, we presented a novel framework for unsupervised anomaly detection in US images using Synomaly noise and a multi-stage diffusion process. Synomaly noise was designed to add synthetic anomalies to healthy US images during training, enabling the model to effectively remove anomalies during inference. The multi-stage diffusion process was introduced to incrementally denoise the images, thereby preserving fine details and generating more realistic reconstructions.
@@
The experiments on carotid US, brain MRI, and liver CT datasets demonstrated the efficacy of the proposed approach. For the carotid US dataset, the proposed method achieved performance comparable to fully supervised segmentation models, demonstrating its strong anomaly removal capabilities even without large amounts of labeled data. The proposed method improved unsupervised anomaly segmentation results by 0.23 in the carotid US dataset. The brain MRI experiments showed an increase of approximately 0.1 in Dice score with the proposed approach, while the liver CT dataset demonstrated an increase of 0.04, showing improved unsupervised anomaly segmentation accuracy. Additionally, the ablation studies highlighted the importance of configuring reasonable hyperparameters for Synomaly noise, as its performance is closely tied to its ability to resemble the target anomalies. We also demonstrated the effectiveness of the multi-stage diffusion process and the masked fusion technique in improving segmentation quality, both of which contributed to enhanced model performance.
@@
Overall, the proposed approach outperformed existing state-of-the-art methods for unsupervised anomaly detection, significantly narrowing the gap between unsupervised and fully supervised approaches. Despite challenges such as variability in acquisition settings and machine types, the proposed method exhibited robustness, generalizability, and promising potential for clinical applications. Moreover, the proposed framework provides an opportunity for doctors to incorporate their prior knowledge of the disease’s appearance, thus enhancing unsupervised anomaly segmentation accuracy by allowing the model to be guided by expert input. Future work could explore training the model to be sensitive to specific types of anomalies while disregarding others, enabling selective segmentation. Such advancements would allow the model not only to detect anomalies but also to classify them to some extent, thereby adding greater value to clinical decision-making and facilitating more tailored medical interventions.
@@
The proposed approach is particularly well-suited for addressing anomalies that are distinguishable based on intensity differences. Anomalies that present clear intensity variations compared to surrounding tissues are efficiently handled by this framework, resulting in high segmentation accuracy. However, the approach is less effective when dealing with anomalies that do not have distinctive intensity contrasts but instead differ in shape or structural characteristics. In such cases, the proposed model does not show evident superiority to differentiate between normal and abnormal shapes, as the generative process is primarily guided by intensity-based information. Future work could explore combining intensity-based anomaly detection with additional spatial or shape features to extend the model’s applicability to a broader range of complex anomalies.

### v2

@@ 4.6.1 Ablation for Multi-stage Diffusion Process
To validate the effectiveness of the proposed multi-stage diffusion process, we conduct an ablation experiment to compare the anomaly segmentation performance with and without multi-stage diffusion. The DDPM was trained using Synomaly noise, and during inference, we tested the model’s performance with a single-stage diffusion approach. To ensure a fair comparison, the optimal inference hyperparameters for single-stage diffusion with Synomaly noise were determined using a grid search and are reported in Tab. 1, denoted as “Synomaly+Single-stage”. In general, the number of noise steps at each stage for multi-stage inference is much smaller than the number of noise steps for single stage, which helps to preserve finer details. The quantitative ablation results can be found in Tab. 3, and the visual results can be found in Fig. 6. It is notable that the reconstructed images shown in Fig. 6 are , used for anomaly mask generation. Due to the generative nature of diffusion, they may exhibit gradual intensity changes even in undetected anomaly regions. However, in the multi-stage framework, the input to the next stage is , where background regions are restored from the original image to prevent over-modification.
@@
| Brain MRI BraTS23 | Liver CT LiTS | Carotid Plaque US | |||||||
| DSC | Precision | Recall | DSC | Precision | Recall | DSC | Precision | Recall | |
| Stage 1 | |||||||||
| Stage 2 | |||||||||
| Stage 3 | |||||||||
| Stage 4 | |||||||||
| Stage 5 | |||||||||
@@
For US carotid plaque anomaly segmentation, implementing the multi-stage diffusion process has improved the Dice score by approximately compared to the single-stage approach. The two visual examples in Fig. 6(g) and Fig. 6(h) also demonstrate that a single diffusion process cannot fully erase the anomalous regions. While the plaque in the upper part of the vessel is successfully removed, the larger plaque at the bottom of the vessel remains in the generated image. After applying multi-stage diffusion process, all the plaques are fully removed. Fig. 6(c) shows the generated images at each stage, clearly illustrating how the plaques are iteratively removed from the original image. Generally, distinct and large anomalies require more iterations to be completely erased. Fig. 6(d) showcases another example involving five iterations, where the plaque at the bottom of the vessel lumen is progressively removed over five iterations, while the smaller plaque at the top is mostly removed in four iterations.
The benefit of the multi-stage diffusion process is most evident in carotid US, where anomalies are more distinct and require stronger corruption to be removed. The multi-stage approach, combined with masked fusion, helps suppress these anomalies while preserving healthy regions. In contrast, for brain MRI and liver CT, where small noise levels are already sufficient to blur the anomaly, the improvement is less significant.
To further evaluate the effectiveness of the proposed multi-stage diffusion process, we calculate the average DSC, precision, and recall of the intermediate anomaly masks at each stage, as summarized in Tab. 4. For this analysis, we include only those test images that underwent five diffusion stages during inference. As shown in Tab. 4, across the multi-stage diffusion process, recall consistently increases, while precision gradually decreases, and DSC improves and eventually stabilizes. This indicates that the model progressively captures more of the anomaly regions over successive stages, resulting in higher recall. The decrease in precision reflects a broader predicted mask that includes more false positives, but the gain in recall outweighs this, leading to improved DSC. By stage four and five, all metrics begin to converge, suggesting that the refinement process reaches a stable segmentation output.
@@
Lastly, we calculated the computational time for both single- and multi-stage diffusion processes. The proposed multi-stage diffusion process introduces additional computational steps compared to single-stage diffusion, resulting in an inference time of 67.65.7 milliseconds per image, compared to 13.85.6 milliseconds for the single-stage approach. The inference speed for other diffusion-based methods with single-stage diffusion is similar to our single-stage implementation. Notably, 67.6 milliseconds per image corresponds to approximately 15 frames per second, which is sufficient for most real-time applications. This ensures that the method remains computationally feasible for practical use while enabling effective anomaly detection.
4.6.2 Ablation for Synomaly Noise
To evaluate the effectiveness of Synomaly noise, we compare the performance of different noise functions when only a single-stage diffusion process is applied. The noise functions considered include Gaussian, Coarse, Pyramid, and Simplex noise, with the corresponding results presented in Tab. 2 as DDPM+Gaussian, DDPM+Coarse, DDPM+Pyramid, and DDPM+Simplex, respectively. The performance of Synomaly noise in a single-stage diffusion process is reported in Tab. 3 as Synomaly+Single-stage.
For brain tumor segmentation, Synomaly noise achieved a Dice score of 0.738, outperforming other noise functions by at least 0.1. Similarly, for carotid plaque segmentation, Synomaly noise yielded a Dice score of 0.688, surpassing the best alternative by over 0.3. In liver tumor segmentation, Synomaly noise alone outperformed the best-performing SOTA noise function (Simplex noise) by 0.01 in Dice score, demonstrating its superiority across all datasets. While the performance gain in liver CT was less pronounced compared to the other datasets, Synomaly noise still consistently outperformed existing noise functions. These results highlight the advantages of Synomaly noise in generating realistic counterfactual images and enhancing anomaly segmentation performance.
4.6.3 Hyperparameter Ablation for Synomaly Noise
In this ablation study, we examine the influence of the hyperparameters of Synomaly noise, as described in Sec. 3.2. The performance of Synomaly noise is closely linked to its ability to represent the target anomalies. If the Synomaly noise generated during training resembles the real abnormalities encountered during inference, the trained model can more effectively erase the diseased areas during testing. The morphology of the generated Synomaly noise is controlled by two parameters: the blurring variance () and the threshold value ().
To evaluate the influence of these parameters, several models were trained with different values of and on the brain MRI, liver CT, and carotid US dataset, respectively. The generated noise is normalized to the range [0, 255], and consequently, the threshold values are also within this range. For brain MRI and liver CT, the threshold value () was fixed at 175, while the blurring variance () varied from 1 to 11. For carotid US, was fixed at 150, with also ranging from 1 to 11. The results are shown in Fig. 7(a), where the red, green, and blue lines represent brain MRI, liver CT, and carotid US, respectively. It is evident that the optimal performance is achieved at for brain MRI and liver CT, whereas for carotid US, the best choice is . Similarly, for , we conducted experiments by fixing at 3 while varying from 90 to 230 for MRI and CT data. For US, was fixed at 7. The results, as shown in Fig. 7(b), indicate that the best performance is achieved when setting to 175 for MRI and CT, while is the best choice for carotid US. Some examples of the generated Synomaly noise are displayed below the plots. The value of controls the size of the synthetic anomaly regions, while determines the number of generated artificial anomaly areas. A smaller value results in smaller synthetic anomaly regions, whereas a lower value leads to a higher number of generated anomaly areas. However, in some cases, when choosing a smaller value, these generated regions sometimes overlap with each other, leading to fewer but larger anomalous areas. According to this ablation study, the optimal parameters for brain MRI and liver CT data are , , while the best combination for carotid US is , , since these parameter choices can best resemble the morphologies of corresponding anomalies based on our experiments.
To facilitate hyperparameter selection without an exhaustive grid search, we provide practical guidelines based on the relative size of anomalous regions. For a 128128 image, we recommend for small anomalies ( of the image), while for moderate anomalies (), such as brain and liver tumor a setting of is suggested. Intermediate cases () can use as a reasonable starting point. For larger anomalies (), such as carotid US plaques, is a suitable choice. These anchor points serve as a structured reference for hyperparameter selection. A smaller generates smaller synthetic anomaly regions, while a lower increases the number of generated anomaly areas. By following these guidelines, users can refine parameters accordingly, reducing computational overhead while maintaining robust anomaly detection performance. Due to the random size distribution of the generated synthetic anomalies (see Fig. 7 lower half), a model trained with a single parameter setting is not strictly limited to detecting anomalies of a specific size but can still generalize to certain out-of-distribution cases.
5 Conclusions
@@
In this paper, we presented a novel framework for unsupervised anomaly detection in medical images using Synomaly noise and a multi-stage diffusion process. Synomaly noise was designed to add synthetic anomalies to healthy images during training, enabling the model to effectively remove anomalies during inference. The multi-stage diffusion process was introduced to incrementally denoise the images, thereby preserving fine details and generating more realistic reconstructions.
@@
The experiments on carotid US, brain MRI, and liver CT datasets demonstrated the efficacy of the proposed approach. For the carotid US dataset, the proposed method achieved performance comparable to fully supervised segmentation models, demonstrating its strong anomaly removal capabilities even without large amounts of labeled data. The proposed method improved unsupervised anomaly segmentation results by 0.23 in the carotid US dataset. The brain MRI experiments showed an increase of approximately 0.1 in Dice score with the proposed approach, while the liver CT dataset demonstrated an increase of 0.04, showing improved unsupervised anomaly segmentation accuracy. Additionally, the ablation studies highlighted the importance of configuring reasonable hyperparameters for Synomaly noise, as its performance is closely tied to its ability to resemble the target anomalies. Our experiments also showed that Synomaly noise alone improves anomaly segmentation performance compared to other noise functions, indicating its effectiveness in generating counterfactual images. We also evaluated the multi-stage diffusion process and masked fusion technique, which contributed to further improvements in segmentation quality. These findings suggest that both Synomaly noise and multi-stage diffusion play important roles in the overall performance of the proposed method.
@@
Overall, the proposed approach outperformed existing state-of-the-art methods for unsupervised anomaly detection, significantly narrowing the gap between unsupervised and fully supervised approaches. Despite challenges such as variability in acquisition settings and machine types, the proposed method exhibited robustness, generalizability, and promising potential for clinical applications. Moreover, the proposed framework provides an opportunity for doctors to incorporate their prior knowledge of the disease’s appearance, thus enhancing unsupervised anomaly segmentation accuracy by allowing the model to be guided by expert input. In the future, hyperparameter selection can also be automated for non-expert users by incorporating a few segmentation examples. Further research could also focus on could also explore training the model to selectively detect specific types of anomalies while disregarding others, enabling selective segmentation. Such advancements would allow the model not only to detect anomalies but also to classify them to some extent, thereby adding greater value to clinical decision-making and facilitating more tailored medical interventions.
@@
The proposed approach is particularly well-suited for addressing anomalies that are distinguishable based on intensity differences. Anomalies that present clear intensity variations compared to surrounding tissues are efficiently handled by this framework, resulting in high segmentation accuracy. However, the approach is less effective when dealing with anomalies that do not have distinctive intensity contrasts but instead differ in shape or structural characteristics. In such cases, the proposed model does not show evident superiority to differentiate between normal and abnormal shapes, as the generative process is primarily guided by intensity-based information. Recent studies have explored integrating structural priors [10], synthetic perturbations [42], and language-guided representations [63] to enhance robustness in such scenarios. Future work could explore combining intensity-based anomaly detection with additional spatial or shape features to extend the model’s applicability to a broader range of complex anomalies [7].
@@
Statement
During the preparation of this work the author(s) used ChatGPT in order to revise the manuscript for fluent and grammatically correct English. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
Appendix A. Dataset Details
A.1 Brain MRI: BraTS Dataset
The BraTS2023 dataset, from the international Brain Tumor Segmentation Challenge [2, 3, 44], is a widely used benchmark in medical imaging. It includes 3D brain MRI scans from tumor patients with pixel-wise annotations for healthy tissue, GD-enhancing tumor (ET), edematous/invaded tissue (ED), and necrotic tumor core (NCR). In this work, all tumor-related labels (ET, ED, NCR) are combined into a single “anomalous” class, yielding two categories: healthy and anomalous tissue.
Since the focus is on tumor detection in 2D images, only axial brain slices are considered in this work and the 2D tumor ground truth segmentation is extracted from the annotated 3D classification volumes. Given that ET, ED and NCR are most commonly found in the cerebral hemispheres, the brain MRI dataset is further refined to include only the central slices. For anomaly detection, the FLAIR sequence is used. All selected slices are resized to pixels to have unified dimensions. anomalous slices are randomly selected as the test set. In total, the dataset consists of healthy training samples, anomalous test samples, and healthy test samples. For normalization, we first clip the intensity values at the 99th percentile to mitigate the effect of extreme outliers, and then apply min-max normalization. To help intuitive understanding, a few representative samples are visualized in Fig. A1(b).
A.2 Liver CT: LiTS Dataset
The Liver Tumor Segmentation Benchmark (LiTS) dataset [13] includes abdominal CT volumes, each with annotations of the liver and tumors. For the purposes of this work, all tumors are categorized as “anomalous” class.
Each 3D abdominal CT volume contains varying axial slices, ranging from and . To prepare the data for model training and evaluation, the annotated segmentations are used to classify slices into healthy and anomalous cases. Further refinement is performed using size thresholding to isolate the relevant slices and all images are resized to pixels. In total, the CT dataset has healthy training images, anomalous test images, and healthy test images. To normalize the images, we first clip the intensity values at the 99th percentile to reduce the impact of extreme outliers, followed by min-max scaling to the [0,1] range. A set of representative samples is shown in Fig. A1(c). These examples highlight the complexity of the anomaly detection task, as the abdominal region exhibits various appearances and the liver itself can differ in size, shape and location. The relatively small ratio between the anomaly and the chest volume can further cause class imbalance problems. These factors make it particularly challenging to accurately segment liver tumors, while preserving healthy tissue. This high diversity and complexity of the LiTS dataset make it a great benchmark for assessing the performance and generalizability of the proposed anomaly detection approaches.
A.3 Carotid US Dataset
The carotid artery US dataset includes US scans from healthy patients and patients with carotid plaques. The US images in the dataset are all captured from the cross-sectional view of the carotid artery. The acquisition was performed within the Institutional Review Board Approval by the Ethical Commission of the Technical University of Munich (reference number 2025-60-S-CB). Due to the nature of US imaging, B-mode appearance is affected significantly by contact point, force, angle, and US machine setting. So, to alleviate the effects caused by the changing image appearance, we crop the US image around the target anatomy, i.e., the carotid artery, to ensure the network only focuses on disease-cased image variations. In this study, we applied existing vessel tracking and segmentation methods to generate vessel masks, which are then used to crop the original full-view US images. Fig. A1(a) presents the pipeline for vessel tracking and cropping procedure. To account for potential imperfections in vessel segmentation and to enhance model robustness, a small degree of randomness regarding the position and size of the vessel within the image is introduced in the cropping process, which also creates some variability in the training data.
The ground truth annotations for plaques in the anomalous carotid artery samples are manually annotated under the supervision of our clinical experts specialized in vascular disease. All such annotations of vessels and anomalies were performed using the ImFusion Suite222https://www.imfusion.com/products/imfusion-suite. All US images are resized to pixels to standardize input dimensions. A min-max normalization is applied to normalize the images to [0,1]. In total, the final dataset contains healthy training images, anomalous test images, and healthy test images. The examples can be found in Fig. A1(d).
Appendix B. Additional Visual Results
Fig. B1, B2, and B3 present additional visual results on some challenging cases in the three datasets.
Fig. B1 shows three extra visual results from the brain MRI dataset. The first two examples illustrate failure cases of the proposed method. In the first, the model incorrectly identifies a small bright region in the upper part of the image as an anomaly. Increasing the value and decreasing the value may improve robustness by making the model less sensitive to small artifacts and more focused on larger anomalous regions. In the second case, the failure is primarily due to the out-of-distribution appearance of the input, which deviates significantly from the training distribution. Incorporating a broader variety of training samples may help mitigate such issues. Nevertheless, the proposed method still demonstrates superior performance compared to other SOTA approaches, even in these challenging cases.
Fig. B2 presents three additional examples from the liver CT dataset. In the first example, the model fails to remove three lesions, likely mistaking them for small hepatic ducts, resulting in missed anomaly detection. Decreasing and increasing during training could improve sensitivity to such subtle lesions. In the second case, the gastrointestinal duct appears as an irregular structure in the lower left but is correctly reconstructed and not mistaken as an anomaly. However, a dark region in the liver is wrongly identified, causing false positives. The third example shows good localization, with most anomalies accurately detected. Despite these challenges, the proposed method consistently outperforms other SOTA approaches.
Fig. B3 presents three additional examples from the carotid US dataset. In the first image, an artifact within the carotid lumen is incorrectly identified as an anomaly and removed in the generated counterfactual image, resulting in a false positive prediction. In the second case, the plaque exhibits a highly distinct boundary and appearance. While the proposed method successfully detects some smaller plaques in the upper region, the larger and more prominent plaque at the bottom is not fully removed. A similar issue is observed in the third example, where the distinct appearance of the plaque hinders its complete removal. To improve the model’s performance on such distinct anomalies, increasing the intensity offset during training can produce more distinct Synomaly noise, enhancing the model’s ability to detect and remove pronounced anomalies. Despite these challenges, the proposed method still demonstrates robust performance compared to other SOTA approaches.